OpenStudy (jolo_yolo):

I need help and im Lost 2. A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 1. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. 2. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). 3. Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

OpenStudy (jolo_yolo):

@ganeshie8 @phi

OpenStudy (phi):

multiplying the first equation by an integer other than 1 if you have an equation like x + 2y= 7 and you multiply it by , for example, 7 that means multiply both sides by 7: 7(x+2y) = 7*7 which involves multiplying *every* term on both sides you would get 7x+14y=49 can you pick a number bigger than 1 ?

OpenStudy (jolo_yolo):

8 - can i use that?

OpenStudy (phi):

yes, any number you like. now multiply 8 times the first equation 2x + 7y = 1

OpenStudy (jolo_yolo):

16x + 56y = 256 ?

OpenStudy (phi):

ok, on the left side. the right side is 8*1

OpenStudy (phi):

8*(2x + 7y ) = 8*1

OpenStudy (phi):

and that is 16x +56y = 8

OpenStudy (phi):

we did this part: multiplying the first equation by an integer other than 1 ok so far? now we do this part: adding it to the second equation. that means add 16x +56y = 8 and -3x – 4y = 5

OpenStudy (phi):

16x+ 56y =8 -3x - 4y = 5

OpenStudy (phi):

add the "x's" 16 x take away 3 x's , gives how many x's do the same for y's and then the 8 + 5

OpenStudy (jolo_yolo):

wait so i take out the variables and add 8 by 5?

OpenStudy (phi):

we add the 2 equations 16x+ 56y =8 -3x - 4y = 5 can you add them ?

OpenStudy (phi):

you add the Left side and you add the right sides

OpenStudy (jolo_yolo):

okay ill try

OpenStudy (jolo_yolo):

sorry my page was disconnected. would it be 13 + 60y = 13 ?

OpenStudy (phi):

16x+ 56y =8 -3x - 4y = 5 now added together: 16x-3x + 56y - 4y = 8+5 can you simplify that ?

OpenStudy (phi):

you are close, but not accurate ... to solve these we have to be careful with the signs

OpenStudy (jolo_yolo):

19x + 52y = 13 ..right?

OpenStudy (phi):

16x-3x is not 19x . the rest is correct

OpenStudy (jolo_yolo):

ayee so 13x

OpenStudy (phi):

yes. so the "new" first equation is 13x+52y= 13 and the second equation is still -3x-4y= 5

OpenStudy (jolo_yolo):

oh okay ty, can u help with part 2 n 3?

OpenStudy (phi):

to solve 13x+52y= 13 -3x-4y= 5 I would first simplify the first equation. Do you see a number that divides evenly into 13, 52 and 13 ?

OpenStudy (jolo_yolo):

52 divided by 2 is 26 but nothing for 13

OpenStudy (phi):

what is 52/4 ?

OpenStudy (jolo_yolo):

13 .~.

OpenStudy (phi):

so the idea is multiply the first equation by 1/13 (i.e. divide each term by 13). can you do that ?

OpenStudy (jolo_yolo):

so 1/3 * 13x + 52y = 13 ?

OpenStudy (phi):

1/13 not 1/3 and we do that to *both sides* and *all terms*

OpenStudy (jolo_yolo):

sorry it was a typo. and ill try to do it

OpenStudy (jolo_yolo):

13x + 52y x 1/13 = 5.30 || 13 x 1/13 = 0.99

OpenStudy (phi):

like this: \[ \frac{1}{13} \cdot 13x +\frac{1}{13} \cdot 52y = \frac{1}{13} \cdot 13 \] think: 13 divided into the "number" part

OpenStudy (phi):

13 x / 13 is x (because 13 divided by itself is 1)

OpenStudy (phi):

what is 52 y / 13 ?

OpenStudy (jolo_yolo):

4y?

OpenStudy (phi):

yes

OpenStudy (phi):

can you write down the simplified version \[ \frac{1}{13} \cdot 13x +\frac{1}{13} \cdot 52y = \frac{1}{13} \cdot 13 \]?

OpenStudy (phi):

the first term is 13 x divided by 13 the second term is 52 y divided by 13 the right hand side (of the = sign) is 13 divided by itself

OpenStudy (jolo_yolo):

i got 1x + 4y = 1

OpenStudy (phi):

yes. and most times we don't bother to write 1x, we write x (but 1x and x mean the samething) to review, we are solving 13x+52y= 13 -3x-4y= 5 we simplified the first equation (by multiplying it by 1/13). we now have 1x + 4y = 1 -3x -4y= 5 what do we get if we add these equations together (be careful of the + and - signs) ?

OpenStudy (jolo_yolo):

-2x -4y = 6

OpenStudy (phi):

1x -3x +4y -4y =1+5 -2x + ? = 6 you did the y part wrong. but the rest is ok. can you re-do the y 's ?

OpenStudy (phi):

4 y's take away 4 y's

OpenStudy (jolo_yolo):

4 + -4 = 0y

OpenStudy (phi):

yes, that means when we add these equations: 1x + 4y = 1 -3x -4y= 5 we get -2x +0 = 6 or just -2x = 6 now divide both sides by -2 to "solve for x"

OpenStudy (jolo_yolo):

x = -3?

OpenStudy (phi):

yes now use 1x + 4y = 1 with x=-3 to find y

OpenStudy (phi):

1x+4y= 1 1*-3 + 4y= 1 -3 + 4y= 1 add +3 to both sides

OpenStudy (phi):

4y=4 divide both sides by 4 y= 1

OpenStudy (phi):

for part 3, use x=-3 and y=1 in the original equations 2x + 7y = 1 -3x – 4y = 5 and show that both are true.

OpenStudy (jolo_yolo):

Okay, thanks for all your help! much appreciation~